OpenStudy (laronjaxon):

how would you solve these? 3x + 2y = 7 4x – 3y = –2

OpenStudy (anonymous):

7/ 2y = 3x

OpenStudy (shane_b):

Wow..there's been about 50 of these types of questions today...what are the schools teaching?! I've typed these about too many times today..someone else GO :)

OpenStudy (laronjaxon):

its on my flvs segment exam :/ i dont get credit forit if i theres no work an i dont even remember learning this :/

OpenStudy (aliahmed):

x=1 y=2

OpenStudy (anonymous):

Ive learned solving eqs. with elimination method however not with two multiplying variables in one equation mor elike 2x + 3 or x/5 - 2 sorts.

OpenStudy (shane_b):

It's very simple. Solve for one unknown variable in terms of the other. Once you have that replace it in the other equation and solve for the single unknown variable. Once you have the unknown variable you can plug it in to either of the original equations and solve for the other unknown. Easy peezy :)

OpenStudy (aliahmed):

y=(7-3x)/3 then sub in eq 2 already u have x& y value

OpenStudy (shane_b):

I guess I'll type it out since no one else has yet:\[3x + 2y = 7\]\[4x – 3y = –2\]First, solve for 1 unknown in terms of the other...I'll use the second equation: \[x=\frac{3y-2}{4}\]Now plug that back into the other equation:\[3(\frac{3y-2}{4})+2y=7\]Solve for y to get y=2\]Now plug y=2 into one of the first equations:\[3x+2(2)=7\]\[x=1\]

OpenStudy (shane_b):

So you end up with x=1 and y=2.

OpenStudy (laronjaxon):

Thank you!